# NUDT4MSTAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild

### v2

@@ 1. $\bullet$
    Richness. The dataset should include rich target categories, scenes, and imaging conditions to approximate the real world and reduce data bias. Besides, the dataset needs enough metadata, such as complex data, scene description, and imaging parameters. Researchers can propose new algorithms by combining metadata with SAR image characteristics.
@@ 3. $\bullet$
Therefore, we build and open-source a new large-scale dataset, NUDT4MSTAR, with various categories and imaging conditions. This dataset is carefully labeled with extensive metadata and has an official benchmark based on 7 experimental settings and 15 comparison methods with 2 data formats. This dataset is our first step in building a new SAR target dataset system. In the future, we will expand the objects from vehicles to a wider category list to develop SAR ATR.
@@ 2.2 Review of SAR Target Classification Algorithms
Early target classification relies on hand-crafting features and classifiers, while contemporary methodologies are predominantly anchored in deep learning-based methods. According to their different feature representations, we divided target recognition methods into four distinct paradigms: traditional features, deep learning, feature fusion, and classification with metadata. In an era increasingly dominated by data-driven algorithms, the availability of a rich, standardized, and usable dataset enriched with metadata is becoming increasingly pivotal.
@@
Traditional features. Previous studies have proposed a plethora of valuable features, encompassing geometric structure features [38, 24, 13], electromagnetic-scattering features [70, 42, 102, 50], time-frequency features [77], local descriptors [15, 76], and sparse representations [32]. These features are particularly tailored to the special coherent imaging mechanisms. For instance, SAR-HOG [76] employs a ratio-based gradient computation rather than the traditional differential gradient approach, and Attributed Scattering Centers (ASC) provides a feature set including location, geometry, and polarization information. Notably, extracting electromagnetic-scattering features relies on complex images with phase information, underscoring the necessity of providing comprehensive data formats.
@@
Deep learning. Deep learning algorithms excel at uncovering intricate patterns within datasets, particularly when ample data is available. Nowadays, these algorithms achieve a remarkable nearly 99% accuracy on MSTAR’s basic Standard Operating Condition (SOC) settings, but the Extended Operating Condition (EOC) with robustness and few-shot are more challenging problems. Convolutional Neural Network (CNN) [45, 63] is a common network in SAR target classification. All-convolutional network (A-ConvNet) [9] was proposed to reduce overfitting and had high accuracy under many experiment settings but is not robust to noise. In recent years, Transformer [85, 16] has also been used in SAR target classification. After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [14, 46, 25], attention mechanism [71, 88], domain alignment [55], reinforcement Learning [40, 8], transfer learning [97], meta-learning [22], physical deep learning [11, 90, 37, 59], semi-supervised learning [106, 80], and self-supervised learning [101, 39, 53]. The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets. Nonetheless, the reliance on outdated datasets and the lack of standardized benchmarks are impeding the advancement of SAR target recognition in the 2020s.
@@
Feature fusion. In order to improve the robustness and interpretability of deep learning methods, many feature fusion methods combining traditional and deep features have been proposed. Kechagias et al. [43] discussed the fusion of deep Learning and sparse coding. FEC [103] used deep CNN and ASC features based on complex data with amplitude and phase. Li et al. [48] combined scattering centers (SC) with graph convolutional network. Besides feature fusion methods, ESF [20] applied ASC features to guide deep model training. EFTL [60] proposed using the ASC model as weight initialization of a complex convolutional network. PIHA [35] proposed a hybrid attention mechanism with ASCs and the deep learning framework. LDSF [96] constructed a heterogeneous graph for the first time to fully exploit both the scattering information of the target components and their interactions. However, many SAR target datasets only provide magnitude images, ignoring phase information and quantization issues.
@@
Classification with metadata. Imaging angles and target-related metadata are the two most prevalent categories of ancillary data. MSTAR and SAMPLE datasets furnish the imaging angles for each respective image. RotANet [92] combined predicting the rotational pattern of an image sequence with self-supervised learning. Zhang et al. [104] investigated the problem of optimal image imaging angle selection in few-shot recognition based on the MSTAR dataset. AaDRL [40] explored the SAR active recognition framework to fully exploit target characteristic variations at different azimuth angles based on the SAMPLE dataset. In addition, many multi-view methods [68, 81] were proposed with image sequences from different imaging angles. Since the target types are known, the researchers used 3D models and shape parameters as metadata for fine-grained tasks. For example, SARbake [67] used 3D CAD models and projection relations for semantic segmentation, and DBAE [91] utilized size and structure attributes for zero-shot classification. Consequently, a SAR target dataset enriched with pertinent metadata, encompassing both imaging parameters and detailed target characteristics, becomes an indispensable step in pursuing accurate and robust SAR target recognition.
@@
Experimental settings. Indeed, the MSTAR dataset has more than 50 types of experimental configurations for SOC and EOC, as detailed in [44]. These configurations include image angle variation, target version/configure variation, noise corruption, occlusion, resolution variation, and background clutter variation [44, 55]. In a similar vein, other SAR classification datasets, such as OpenSARShip and FUSAR-Ship, lack a standard experimental benchmark, and researchers also devise different bespoke settings that leverage subsets of the target class for classification [37, 59].
@@
Evaluation metrics. Classification accuracy and confusion matrix are the primary metrics employed for evaluating the performance of classification models. Accuracy quantifies the proportion of correctly classified samples relative to the total number of samples in the dataset. The confusion matrix describes the correspondence between the classification result and the ground truth. Within this matrix, each column corresponds to the predicted class labels, whereas each row represents the true class labels. Besides, the Receiver Operating Characteristic (ROC) curve is a valuable tool for assessing used for binary classification and detection tasks such as suppressing background clutter and outlier rejection [9].
@@
Figure 5: Taxonomic systems of NUDT4MSTAR. For civilian vehicles, the taxonomy is based on Chinese and European vehicle classification standards [82, 19], according to the vehicle’s purpose, structure, size, and mass. For military vehicles, we followed the MSTAR taxonomic system [73]. (a) Taxonomic systems of NUDT4MSTAR. Our dataset is a comprehensive civilian vehicle map covering 4 main categories, 21 classes, and 40 types. We provide a detailed illustration of the histogram distribution for these 40 vehicle types, demonstrating the breadth and depth of our dataset. Its sufficient and balanced samples make it well-equipped to meet various experimental settings and studies. (b) Target classes of SAR vehicle datasets. We statistically analyze the number of civilian and military vehicle classes and types in SAR vehicle datasets. Our dataset greatly enhances the civilian vehicle richness for the SAR ATR, compared to the measured dataset Gotcha and the simulated datasets CVDomes and SARSim. (c) List of vehicle abbreviations.
@@
We noticed a pronounced preference for the MSTAR dataset among SAR classification algorithms. Most algorithms are based on the MSTAR dataset, with a small number using other datasets such as SAMPLE, OpenSARShip, and FUSAR-Ship. This preference is attributed to the distinctive benefits of the MSTAR dataset released in the 1990s. As depicted in Fig. 4 (a), studies requiring complex data have opted for the MSTAR dataset, whereas other datasets that only provide magnitude data cannot be used to validate their algorithms. In addition, information about vehicle types in the dataset is also available on the web. Compared to other datasets with long-tailed distributions, the MSTAR dataset has a more balanced target sample and diverse imaging conditions, which facilitates the establishment of experimental frameworks and the investigation of recognition challenges. Therefore, MSTAR has accumulated a large number of algorithms and experimental settings. However, this dataset has not released all target samples, and the extant algorithms have achieved peak performance in prevalent settings, such as classification with ten target types, imaging angle variation, and version variation. The ideal acquisition conditions of the MSTAR dataset cannot show that the recognition problem is truly well solved. Scholars are endeavoring to address these limitations by proposing novel experimental settings and datasets. However, as shown in Fig. 4 (b), some simulation experimental settings based on measured data may not accurately reflect the real world. Besides, the sample diversity is difficult to ensure due to collection and labeling costs. Therefore, we perceive a constraint on advancing SAR ATR techniques due to data limitations. Constructing a new open-source dataset benchmark with diverse target categories and imaging conditions is the cornerstone of excellent and stable SAR ATR techniques.
3 Building NUDT4MSTAR: Why, What, How
@@
In this section, we present NUDT4MSTAR from three perspectives: (1) Why we constructed this dataset in the 2020s; (2) What we did to ensure its diversity and standardization; (3) How the diversity and statistical analysis of this dataset to support SAR ATR.
@@
TABLE III: Statistics of NUDT4MSTAR. We segment the targets with fixed-size slices and random offsets to investigate their characteristics and aim for robust recognition under different collection conditions. We varied the target location in urban and factory scenes and collected data at the factory several times. In addition, some targets can not be labeled due to heavy occlusion. # Types.: Number of object types. Res.: Resolution. Pol.: Polarization. Dep.: Depression angle. Azi.: Target azimuth angle interval. Img. Size: image size. # Img.: Number of images in ground and slant range coordinate systems.
Scene # Types Platform Mode Band Res. (m) Pol. Dep. (°) Azi. (°) Img. Size # Img.
[…]
@@
The emergence of new datasets and papers has led to rapid progress in various areas of SAR ATR, demonstrating the attention given to this issue in recent years. However, a large-scale dataset with various target categories and imaging conditions is still lacking because of SAR’s expensive costs and annotation. We aim to achieve as much diversity as possible in target categories, scene settings, and sensor conditions. With this in mind, our dataset has huge potential to advance the field of SAR ATR significantly. Its values for SAR ATR are:
@@ 3.1 Construct program
Target. First of all, according to the Chinese and European vehicle classification standards [82, 19], we confirm the main target categories based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special). Then, we subdivide each category into various target classes based on size and special structure and collect as many different types as possible. Finally, NUDT4MSTAR has 4 categories, 21 classes, and 40 types. It covers most civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets.
Figure 7: Demo images, classes, types, and sizes of forty vehicles in NUDT4MSTAR. Here are the SAR (pseudo-color for better visualization of SAR magnitude images) and corresponding RGB images. As we can see, the obvious difference between the vehicles in the SAR images is the scattering characteristic variation due to size and structure. Therefore, we classify various vehicle types based on their structure and size. Their length, width, and height measurements are listed in meters (m). Besides, it is worth noting that our SAR images are not in the ideal situation where the target is located right in the center of the demo image, just as in the MSTAR dataset, but has a random offset similar to the QinetiQ dataset.
@@
Table III shows the open-source data list. After a year and a half’s effort (from April 2023 to October 2024), we completed program designation, data acquisition and processing, and benchmark construction. We commissioned the Institute of Stealthy Engineering Technology Company in Beijing to complete the data collection and annotation.
@@ 3.2 Data acquisition
Annotation. We annotate the target types and positions with rectangular box labels according to the optical images and placement positions. Since we focus on individual target characteristics, the targets are placed at a certain distance from each other. Besides our vehicles, other vehicles in the scene are labeled as “other”. After labeling, we acquired the target slices and added a random offset.
@@ 3.3 Professional system
[…]
@@ 3.4 Statistical Analysis
We now discuss the properties of our dataset and provide some statistics compared with others.
Class distribution. NUDT4MSTAR includes 4 categories, 21 classes and 40 types of vehicles with balanced samples, as shown in Fig. 5. Relative to extant SAR target datasets, our dataset’s targets’ enhanced diversity and intricacy introduce novel formidable challenges to the SAR target fine-grained recognition research. Moreover, adequate samples guarantee the availability of a substantial array of target classes across experimental settings, facilitating more diverse investigations.
@@
Other vehicle targets. Other vehicle targets in the scene are also labeled for the potential detection task in large remote images. However, this dataset mainly discusses the scattering properties of individual targets. These interfering vehicles can be used as part of a future rejection issue.
@@
Figure 10: Influence of different scenes on target characteristics in SAR images. Most current SAR target classification datasets are collected in simple scenes, e.g., (a) MSTAR is collected in a flat grass scene with a clean background and pseudo-correlation between target and background. While (b) our NUDT4MSTAR samples are collected under different positions in various scenes. In Fig. (b), it is clear that the same target under the same imaging angle has obvious signature changes due to different scenes. For example, other objects higher up in front of and behind the target can create occlusions and layovers, reducing reflected energy and increasing non-target scattering. Besides, the target shadow in difficult scenes is not as obvious as the MSTAR data. (c) Occlusion and layover. Compared to the bare soil scene without interfering objects, the factory and woodland scenarios indicate occlusion and layover. Shadows from interfering objects in the former may obscure targets, while trees in the latter are likely to increase non-target scattering. Due to occlusion and layover are complex results of a combination of target, interference, and imaging geometry, we illustrate this problem with a single target demo. These statistics are from an SUV Chang’an CSCS75 Plus (Vehicle 12) at different image angles in the 3 scenes’s ground range images.
Occlusion in scenes. Previous vehicle datasets focus on target characteristics with simple scenes, and we collect samples from different scenes with the same target and angle. Fig. 10 shows their different influences with occlusions and layovers in different scene regions. The shadow of buildings and roadside trees in the factory may obscure targets and reduce target scattering, and the nearby trees in woodland have more obvious occlusion and increase non-target scattering.
Figure 11: Target azimuth angle distribution of (a) MSTAR and (b) NUDT4MSTAR. Compared to the MSTAR dataset, which has depression angles mainly publicized at 17° and 15° with incomplete angles for most target classes, our NUDT4MSTAR dataset provides balance and comprehensive angles for all targets. However, the sampling interval of target angles is sparser due to strip imaging.
Imaging angle. Different imaging angles affect the target scattering characteristics, and the imaging geometry relationship can change the interference in the scene. Our dataset has various imaging angles with balance distribution in different depression angles, as shown in Fig. 11. However, the sampling interval of target azimuth angles is not equally dense in different scenes due to the limitation of stripmap imaging and cost problems.
@@
Quality control. To ensure high-quality data, we considered challenges such as target and site hire, airspace applications, weather conditions, and data checking. Eight people completed the data acquisition over six months. In addition, 14 labelers and two inspectors perform the data annotation about four months.

### v3

@@ 1. $\bullet$
    Richness. A comprehensive dataset should include diverse target categories, scenes, and imaging conditions to reflect real-world environments and reduce biases. It also needs to contain enough formats and metadata, such as complex data and imaging parameters, to help researchers in algorithm innovation.
@@ 3. $\bullet$
Therefore, we build and open-source a new large-scale dataset, ATRNet-STAR, with various categories and imaging conditions. This dataset is carefully labeled with extensive metadata and has an official benchmark based on 7 experimental settings and 15 comparison methods with 2 data formats. This dataset is our first step in building a new SAR target dataset system. In the future, we will expand the objects from vehicles to a wider category list to further promote SAR ATR research.
@@ 2.2 Review of SAR Target Classification Algorithms
Early target classification relies on hand-crafting features and classifiers, while contemporary methodologies are predominantly anchored in deep learning-based methods. According to their different feature representations, we divided target recognition methods into four distinct paradigms: traditional features, deep learning, feature fusion, and classification with metadata. In an era increasingly dominated by data-driven algorithms, the availability of a rich, standardized, and usable dataset is becoming increasingly pivotal.
@@
Traditional features. Previous studies have proposed a wide range of valuable features, encompassing geometric structure features [54], electromagnetic-scattering features [55, 56, 57, 58], time-frequency features [59], local descriptors [7, 60], and sparse representations [61]. These features are particularly tailored to the special coherent imaging mechanisms. For instance, SAR-HOG [60] employs a ratio-based gradient computation rather than the traditional differential gradient approach, and Attributed Scattering Centers (ASC) provides a feature set including location, geometry, and polarization information. Notably, extracting electromagnetic-scattering features relies on complex images with phase information, underscoring the necessity of providing comprehensive data formats.
@@
Deep learning. Deep learning algorithms excel at uncovering intricate patterns within datasets, particularly when ample data is available. Nowadays, these algorithms achieve remarkable accuracy, reaching nearly 99% on MSTAR’s basic Standard Operating Condition (SOC) settings. However, the Extended Operating Condition (EOC), which includes robustness and few-shot, remains a significant challenge. Convolutional Neural Network (CNN) [62, 63, 64] is a common network in SAR target classification. All-convolutional network (A-ConvNet) [65] was proposed to reduce overfitting and had high accuracy under many experiment settings but is not robust to noise. In recent years, Transformer [66, 67, 68] emerged as a powerful alternative, offering new perspectives on feature extraction. After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [69, 70], attention mechanism [71, 72], domain alignment [53], reinforcement Learning [73, 74], transfer learning [75], meta-learning [76, 77, 78], physical deep learning [79, 80, 81], semi-supervised learning [82], and self-supervised learning [83, 29]. The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets. Nonetheless, the reliance on outdated datasets and the lack of standardized benchmarks are impeding the advancement of SAR target recognition in the 2020s.
@@
Feature fusion. In order to improve the robustness and interpretability of deep learning methods, many feature fusion methods combining traditional and deep features have been proposed. Kechagias et al. [84] discussed the fusion of deep learning and sparse coding. FEC [85] used deep CNN combined with ASC features derived from complex data, incorporating both amplitude and phase information. Li et al. [86] combined scattering centers (SC) with graph convolutional network. Besides feature fusion methods, ESF [87] applied ASC features to guide deep model training. EFTL [88] proposed using the ASC model as weight initialization of a complex convolutional network. PIHA [89] proposed a hybrid attention mechanism with ASCs and the deep learning framework. LDSF [90] constructed a heterogeneous graph for the first time to fully exploit both the scattering information of the target components and their interactions. However, many SAR target datasets only provide magnitude images, ignoring phase information and quantization issues.
@@
Classification with metadata. Imaging angles and target-related metadata are the two most prevalent categories of ancillary data. MSTAR and SAMPLE datasets furnish the imaging angles for each respective image. RotANet [91] combined predicting the rotational pattern of an image sequence with self-supervised learning. Zhang et al. [92] investigated the problem of optimal image imaging angle selection in few-shot recognition based on the MSTAR dataset. AaDRL [73] explored the SAR active recognition framework to fully exploit target characteristic variations at different azimuth angles based on the SAMPLE dataset. In addition, multi-view methods [93] were proposed with image sequences from different imaging angles. Since the target types are known, the researchers used 3D models and shape parameters as metadata for fine-grained tasks. For example, SARbake [94] used 3D CAD models and projection relations for semantic segmentation, and DBAE [95] utilized size and structure attributes for zero-shot classification. Consequently, a SAR target dataset enriched with pertinent metadata, encompassing both imaging parameters and detailed target characteristics, becomes an indispensable step in pursuing accurate and robust SAR target recognition.
@@
Experimental settings. Indeed, the MSTAR dataset has more than 50 types of experimental configurations for SOC and EOC, as detailed in [11]. These configurations include image angle variation, target version/configure variation, noise corruption, occlusion, resolution variation, and background clutter variation [11, 53]. In a similar vein, other SAR classification datasets, such as OpenSARShip and FUSAR-Ship, lack a standard experimental benchmark, and researchers also devise different bespoke settings that leverage subsets of the target class for classification [81].
@@
Evaluation metrics. Classification accuracy and confusion matrix are the primary metrics employed for evaluating the performance of classification models. Accuracy quantifies the proportion of correctly classified samples relative to the total number of samples in the dataset. The confusion matrix provides a detailed correspondence between the predicted and true class labels, with each column representing the predicted class and each row representing the true class. Besides, the Receiver Operating Characteristic (ROC) curve is a valuable tool for assessing used for binary classification and detection tasks such as suppressing background clutter and outlier rejection [65].
@@
Figure 5: Taxonomic systems of ATRNet-STAR. For civilian vehicles, the taxonomy is based on Chinese and European vehicle classification standards [96, 97], according to the vehicle’s purpose, structure, size, and mass. For military vehicles, we followed the MSTAR taxonomic system [98]. (a) Taxonomic systems of ATRNet-STAR. Our dataset is a comprehensive civilian vehicle map covering 4 classes, 21 subclasses, and 40 types. We provide a detailed illustration of the histogram distribution for these 40 vehicle types, demonstrating the breadth and depth of our dataset. Its sufficient and balanced samples make it well-equipped to meet various experimental settings and studies. (b) Target classes of SAR vehicle datasets. We statistically analyze the number of civilian and military vehicle classes and types in SAR vehicle datasets. Our dataset greatly enhances the civilian vehicle richness for the SAR ATR, compared to the measured dataset Gotcha and the simulated datasets CVDomes and SARSim. (c) List of vehicle abbreviations.
@@
We noticed a pronounced preference for the MSTAR dataset among SAR classification algorithms. Most algorithms are based on the MSTAR dataset, with a small number using other datasets such as SAMPLE, OpenSARShip, and FUSAR-Ship. This preference is attributed to the distinctive benefits of the MSTAR dataset released in the 1990s. As depicted in Fig. 4 (a), studies requiring complex data have opted for the MSTAR dataset, whereas other datasets that only provide magnitude data cannot be used to validate their algorithms. In addition, information about vehicle types in the dataset is also available on the web. Compared to other datasets with long-tailed distributions, the MSTAR dataset has a more balanced target sample and diverse imaging conditions. Therefore, MSTAR has accumulated a large number of algorithms and experimental settings. However, this dataset has not released all target samples, and the extant algorithms have achieved peak performance in common popular settings [11], such as classification with ten target types, imaging angle variation, and version variation. The ideal acquisition conditions of the MSTAR dataset cannot show that the recognition problem is truly well solved. Scholars aim to address these limitations by proposing novel simulation experimental settings. However, as shown in Fig. 4 (b), some simulation experimental settings may not accurately reflect the real-world conditions. Besides, ensuring sample diversity is fraught with challenges due to the high data collection and labeling costs. These limitations collectively impose significant constraints on the advancement of SAR ATR techniques. Constructing a new open-source dataset benchmark with diverse target categories and imaging conditions is a challenging cornerstone of excellent and stable SAR ATR techniques.
3 Building ATRNet-STAR: Why, What, How
@@
In this section, we present ATRNet-STAR from three perspectives: (1) Why we constructed this dataset in the 2020s; (2) What we did to ensure its diversity and standardization; (3) How the diversity and statistical analysis of this dataset to support SAR ATR.
@@
TABLE III: Statistics of ATRNet-STAR. We segment the targets with fixed-size slices and random offsets to investigate their characteristics and aim for robust recognition under different collection conditions. We varied the target location in urban and factory scenes and collected data at the factory several times. In addition, some targets can not be labeled due to heavy occlusion. # Types.: Number of object types. Res.: Resolution. Pol.: Polarization. Dep.: Depression angle. Azi.: Target azimuth angle interval. Img. Size: image size. # Img.: Number of images in ground and slant range coordinate systems.
Scene # Types Platform Mode Band Res. (m) Pol. Dep. () Azi. () Img. Size # Img.
City 40 airborne strimap X 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 5 128 83,465
Factory 40 airborne strimap X & Ku 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 30 128 63,597
Sandstone 40 airborne strimap X & Ku 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 30 128 30,720
Woodland 11 airborne strimap X & Ku 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 30 128 8,094
Bare soil 11 airborne strimap X & Ku 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 30 128 8,448
@@
The emergence of new datasets and papers has led to rapid progress in various areas of SAR ATR, demonstrating the attention given to this issue in recent years. However, a large-scale dataset with various target categories and imaging conditions is still lacking because of SAR’s expensive costs and annotation. We aim to achieve as much diversity as possible in target categories, scene settings, and sensor conditions. Such multidimensional design establishes it as an instrumental resource for advancing SAR ATR research. Its principal values for the SAR ATR are fourfold:
@@ 3.1 Construct program
Target. First of all, according to the Chinese and European vehicle classification standards [96, 97], we confirm the main target classes based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special). Then, we subdivide each class into various subclasses based on size and special structure and collect as many different types as possible. Finally, ATRNet-STAR has 4 classes, 21 subclasses, and 40 types. It covers most civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets.
Figure 7: Demo images, classes, types, and sizes of forty vehicles in ATRNet-STAR. Here are the SAR (pseudo-color for better visualization of SAR magnitude images) and corresponding RGB images. As we can see, the obvious difference between the vehicles in the SAR images is the scattering characteristic variation due to size and structure. Therefore, we classify various vehicle types based on their structure and size. Their size (length * width * height) measurements are listed in meters (m). Besides, it is worth noting that our SAR images are not in the ideal situation where the target is located right in the center of the demo image, just as in the MSTAR dataset, but has a random offset similar to the QinetiQ dataset.
@@
Table III presents the open-source data list. After a year and a half’s effort (from April 2023 to October 2024), we completed program designation, data acquisition and processing, and benchmark construction. We commissioned the Institute of Stealthy Engineering Technology Company in Beijing to complete the data collection and annotation.
@@ 3.2 Data acquisition
Annotation. Target categories and coordinates are annotated using rectangular box labels based on optical reference images and deployment records. Focusing on individual target signatures, all objects maintain specified separation distances during placement. Besides our vehicles, other vehicles in the scene are labeled as “other”. After labeling, we acquired the target slices and added a random offset.
@@ 3.3 Professional system
Data format. Due to the unique imaging mechanisms, processing steps, and image properties of SAR, we offer different levels of products to support the comprehensive study of the impact of imaging steps on recognition problems. We provide magnitude images in the ground range coordinate system with 8-bit and 32-bit uint formats. The magnitude images with 16-bit uint and complex data with 32-bit float in the slant range coordinate system are also available. In subsequent experimental benchmarks, we use 8-bit ground distance magnitude images and 32-bit slant distance complex data. We also perform other processing based on the slant distance data, such as phase gradient autofocus and polarization decomposition, but the current version of this dataset does not include the processed data after these steps. Raw echoes are also stored, but due to the large storage ($\sim$25TB), they will not be available for upload and download. Researchers can contact us to obtain raw echo if needed. We hope these products will help the researchers easily work with our dataset while providing more tractability to explore and discuss specific imaging mechanisms and image properties for SAR recognition applications.
@@ 3.4 Statistical Analysis
We present key characteristics and comparative statistics of our dataset relative to others.
Class distribution. ATRNet-STAR includes 4 class, 21 subclasses and 40 types of vehicles with balanced samples, as shown in Fig. 5. Relative to extant SAR target datasets, our dataset’s targets’ enhanced diversity and intricacy introduce novel formidable challenges to the SAR target fine-grained recognition research. Moreover, adequate samples guarantee the availability of a substantial array of target classes across experimental settings, facilitating more diverse investigations.
@@
Figure 10: Influence of different scenes on target characteristics in SAR images. Most current SAR target classification datasets are collected in simple scenes, e.g., , (a) MSTAR is collected in a flat grass scene with a clean background and pseudo-correlation between target and background. While (b) our ATRNet-STAR samples are collected under different positions in various scenes. In Fig. (b), it is clear that the same target under the same imaging angle has obvious signature changes due to different scenes. For example, other objects higher up in front of and behind the target can create occlusions and layovers, reducing reflected energy and increasing non-target scattering. Besides, the target shadow in difficult scenes is not as obvious as the MSTAR data. (c) Occlusion and layover. Compared to the bare soil scene without interfering objects, the factory and woodland scenarios indicate occlusion and layover. Shadows from interfering objects in the former may obscure targets, while trees in the latter are likely to increase non-target scattering. We illustrate this problem with a single target demo at the same angle across scenes, but occlusion and layover result from a combination of target, interference, and imaging geometry. These statistics are from an SUV Chang’an CSCS75 Plus (Vehicle 12) at different image angles in the 3 scenes’s ground range images.
Occlusion in scenes. Whereas previous vehicle datasets prioritize target characteristics in ideal environments, we systematically acquire samples across diverse scenes. Fig. 10 shows their different influences with occlusions and layovers in different scene regions. The shadow of buildings and roadside trees in the factory may obscure targets and reduce target scattering, and the nearby trees in woodland have more obvious occlusion and increase non-target scattering.
Figure 11: Target azimuth angle distribution of (a) MSTAR and (b) ATRNet-STAR. Compared to the MSTAR dataset, which has depression angles mainly publicized at 17 and 15 with incomplete angles for most target classes, our ATRNet-STAR dataset provides balance and comprehensive angles for all targets. However, the sampling interval of target angles is sparser due to strip imaging.
Imaging angle. Different imaging angles affect the target scattering characteristics, and the imaging geometry relationship can change the interference in the scene. Our dataset has various imaging angles with balance distribution in different depression angles, as shown in Fig. 11. However, azimuth angle sampling density exhibits scene-dependent due to stripmap mode constraints and operational cost limitations.
@@
Quality control. To ensure high-quality data, we considered challenges such as target and site hire, airspace applications, weather conditions, and data checking. Eight people completed the data acquisition over six months. In addition, 14 labelers and two inspectors perform the data annotation about four months. The entire project cost nearly two years.